OpenStudy (anonymous):

Which of the following is a geometric sequence? -24, 12, 6 -24,23,-22 -24, -12, 0 24, 12, 6

OpenStudy (angela210793):

geometric progression----> yn/y(n-1)=y3/2=y2/y1=q so the forth is a geometric progression and q=1/2

OpenStudy (amistre64):

if a,b,c,d is a geometric sequence then: d c b -- = -- = --; right? c b a